# Entanglement entropy bounds for pure states of rapid decorrelation

### v1

@@ 4.2 Area-law bound
    For $d=2$, under the assumption of frustration freeness of a gapped local Hamiltonian on $W\subset\mathbb{Z}^{2}$, Anshu, Arad and Vidick [9] devised a technique for constructing an approximate ground state projector based on a so-called detectability argument. Using that, Anshu, Arad and Gosset [8] established an area-law bound with a poly-logarithmic correction in case of ground states and $A\subset W$ cylindrical sets.
Beyond this two-dimensional case, only conditional results with assumptions that do not cover the full non-critical regime are available [43, 21, 15], and the area law is known to fail for some non-regular graphs [2].
Since the Rényi entropy $S_{\alpha}(\varrho):=(1-\alpha)^{-1}\ln\mathrm{tr}\varrho^{\alpha}$, which agrees with the von Neumann entropy for $\alpha=1$, is monotone decreasing in $\alpha\in(0,\infty)$, the above area-law bound extends to all values $\alpha\geq 1$. In particular, this applies to integer values $\alpha\in\{2,3,\dots\}$ in $d=1$, for which Cardy and Calabrese [20] proposed a value of the proportionality constants of the area law related to the central charge and the correlation length of the conformal field theory describing the non-critical ground state, whose entanglement entropy is described.
@@ 4.3 Entropy approximations
$\frac{1}{2}\Delta_{l}$ $:=\frac{1}{2}\sum_{j=1}^{\nu^{|B_{l}|}}\left|\lambda_{j}^{\downarrow}-\mu_{j}^{\downarrow}\right|\leq\frac{1}{2}\left\|\varrho_{A}-\widehat{\varrho}_{A}\right\|_{1}$
$\leq\frac{1}{2}\left\||\psi\rangle\langle\psi|-|\psi_{(B_{l})}\rangle\langle\psi_{(B_{l})}|\right\|_{1}=\sqrt{1-\left|\langle\psi|\psi_{(B_{l})}\rangle\right|^{2}}$
@@
where we used [12, Lemma IV.3.2], the variational characterization of the trace norm, $\|a\|_{1}=\sup_{\|b\|=1}\left|\mathrm{tr}(ab)\right|$, as well as the Fuchs-van der Graaf inequality (3.1) in the first three inequalities. The last line is from the fidelity bound (3.7) and the assumed decay. The Fannes-type entropy bound from Proposition A.1 thus yields
@@
In the subsequent analysis of the mutual information, we need a minor modification of Theorem 4.5 involving the approximation of $\psi$ by the vector $|\psi_{(B_{l})}\rangle$ from (2.7) with phases now optimized according to (3.6). For simplicity, we restrict attention to the case $n=2$, i.e., the phases will be adapted to the decomposition of $A$ into two separated components $A_{1},A_{2}\subset W$, cf. Figure 2. To formulate a bound, we require the following modification of Defintion 4.2.
@@
Remark 4.7.
It will be shown in Theorem 6.1 that in the above situation
@@
The proof starts by noting that for any of the three reduced states $\varrho_{D}\equiv\varrho_{D}(\psi)$ with $D\in\{A,A_{1},A_{2}\}$, Uhlmann’s bound (3.2) and the fidelity estimate (3.7) still yield for any $l\geq l_{0}(A)$:
@@ 5 Mutual information bounds
Let $|\psi\rangle$ be a state, which is exponentially decaying at $l_{0}(A)=c\ln|\partial A|$ for some $c\in(0,\infty)$. Let $A_{1},A_{2}\subset W$ be non-empty squares, which are separated by distance $d(A_{1},A_{2})\geq 3l_{0}(A)$. Then there are constants $C,\kappa,\eta\in(0,\infty)$, which are independent of $A_{1},A_{2}$, such that
@@
Moreover, as is evident from the subsequent proof, Theorem 5.1 admits a generalisation to the mutual information of finitely many disjoint squares $A_{1},\dots,A_{n}$.
@@
    12($\frac{1}{2}\left(1-\left|\langle\psi^{(12)}_{(B)}|\psi_{(B)}\rangle\right|^{2}\right)$ $\leq\frac{1}{2}\sum_{\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B}}\Big{|}\ p(\boldsymbol{\sigma}_{B})p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B})-\prod_{j=1}^{2}p(\boldsymbol{\sigma}_{B_{j}})p(\boldsymbol{\sigma}_{A_{j}}|\boldsymbol{\sigma}_{B_{j}})\Big{|}$
@@
The trace distances occuring in the application of Proposition A.1 are estimated by fidelities using (3.1). In turn, each of the fidelities is estimated using Uhlmann’s bound (3.2) in terms of
$1-\left|\langle\psi^{(12)}_{(B)}|\psi_{(B)}\rangle\right|^{2}$ $\leq 2\delta\left(A_{1}\sqcup B_{1}|A_{2}\sqcup B_{2}\right)$
$\leq 2e^{-(l-l0(A))/(3\xi)}=2\ |\partial A|^{c/(3\xi)}e^{-l/(3\xi)}\,.$ (5.6)
Here, the first inequality is (2) and the second follows from the assumption and the fact that $d(A_{1}\sqcup B_{1},A_{2}\sqcup B_{2})\geq l/3$ by construction. Since $|\partial A|\leq 2\max\left\{|\partial A_{1}|,|\partial A_{2}|\right\}$, this completes the proof. ∎
@@ 6.1 General properties
${\delta}(A\,|\,C)=\frac{1}{2}\max_{\|F_{A,C}\|_{\infty}\leq 1}\sum_{\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{C}}F_{A,C}(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{C})\left[p(\boldsymbol{\sigma}_{A}\boldsymbol{\sigma}_{C})-p(\boldsymbol{\sigma}_{A})p(\boldsymbol{\sigma}_{C}))\right]\,,$ (6.2)
where the maximum is over $F_{A_{C}}:\Omega_{A}\times\Omega_{C}\to[-1,1]$ and attained at $F(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{C}):=\operatorname{sign}[p(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{C})-p(\boldsymbol{\sigma}_{A})p(\boldsymbol{\sigma}_{C})\big{]}$. Monotonicity in $C$ follows, since replacing $\boldsymbol{\sigma}_{C}$ by $\boldsymbol{\sigma}_{C\sqcup D}$ only enlarges the collection of functions over which the maximization is performed.
@@
${\delta}_{B}(A,\{u\})\leq\sum_{\boldsymbol{\sigma}_{B}}p(\boldsymbol{\sigma}_{B})\ \overline{{\mathrm{TV}}}_{\boldsymbol{\sigma}_{B}}(A|\{u\})\,.$ (6.6)
@@ 6.2 Relation to correlation function using the FKG property
    and $\langle\sigma_{a};\sigma_{u}\rangle_{\boldsymbol{\sigma}_{B}}:=\langle\sigma_{u}\sigma_{v}\rangle_{\boldsymbol{\sigma}_{B}}-\langle\sigma_{u}\rangle_{\boldsymbol{\sigma}_{B}}\langle\sigma_{v}\rangle_{\boldsymbol{\sigma}_{B}}$ with the subscript indicating that the expectation values are with respect to the conditional probability $p(\cdot|\boldsymbol{\sigma}_{B})$, conditioned on the value of $\boldsymbol{\sigma}_{B}$.
@@
$=\min\left\{\mathbb{P}\left(\boldsymbol{\sigma}_{A}^{(1)}\neq\widetilde{\boldsymbol{\sigma}}_{A}^{(2)}\right)\ \big{|}\ \mbox{couplings with $\boldsymbol{\sigma}_{A}^{(1)}\rightsquigarrow\boldsymbol{\sigma}_{A}^{+},\boldsymbol{\sigma}_{A}^{(2)}\rightsquigarrow\boldsymbol{\sigma}_{A}^{-}$}\right\}\,.$
@@
$\mathbb{P}\big{(}\boldsymbol{\sigma}_{A}^{(1)}\big{)}\neq\boldsymbol{\sigma}_{A}^{(2)}\big{)}$ $\leq\mathbb{E}\left(\sum_{a\in A}(\sigma_{a}^{(1)}-\sigma_{a}^{(2)})/2\right)$
$=\frac{1}{2}\sum_{\boldsymbol{\sigma}_{A}}M_{A}(\boldsymbol{\sigma}_{A})\big{[}p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B}\sigma_{u}=+1)-p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B}\sigma_{u}=+1)\big{]}$
@@
$=\frac{1}{2}\sum_{\sigma_{u},\widehat{\sigma}_{u}}\sum_{\boldsymbol{\sigma}_{A},,\widehat{\boldsymbol{\sigma}}_{A}}p(\boldsymbol{\sigma}_{A},\sigma_{u}|\boldsymbol{\sigma}_{B})\ p(\widehat{\boldsymbol{\sigma}}_{A},\widehat{\sigma}_{u}|\boldsymbol{\sigma}_{B})\left[M_{A}(\boldsymbol{\sigma}_{A})-M_{A}(\widehat{\boldsymbol{\sigma}}_{A})\right]\left[\sigma_{u}-\widehat{\sigma}_{u}\right]$
$=2\,p_{\boldsymbol{\sigma}_{B}}(+)\ p_{\boldsymbol{\sigma}_{B}}(-)\sum_{\boldsymbol{\sigma}_{A}}M_{A}(\boldsymbol{\sigma}_{A})\big{[}p(\boldsymbol{\sigma}_{A}|\sigma_{u}=+1)-p(\boldsymbol{\sigma}_{A}|\sigma_{u}=+1)\big{]}$ (6.13)
@@ 7 Application to the quantum Ising model
    FKG: Like its classical version [38, 32], the quantum Ising model’s probabilty has the FKG property, in the natural partial order of the stoquastic $S^{z}$ basis. This makes Theorem 6.4 applicable to this model, with its variables $\sigma_{u}$ interpreted as $S^{z}_{u}$.
@@
    In its classical version, the inequality is a direct implication of the recently formulated Ising inequality, which was presented in [23, Eq. (1.4)]. The validity of its quantum version follows by a continuity argument which, though not spelled, is indicated above. An alternative direct proof, of both classical and quantum versions using the random current approach, is included in [3].
@@
As noted above, in the $z$-diagonalising computational basis, the ground states of the quantum Ising model (in finite volumes $W\subset\mathbb{Z}^{d}$ with free boundary conditions) are stoquastic states, with the FKG property.
By the extended Ding-Song-Sun inequality (7.4), for any finite volume and throughout the range of the model’s coupling constants, its ground state’s truncated correlation function, conditioned on arbitrary configuration of the buffer, satisfies:
@@
where the last upper bound holds thanks to the Griffiths inequality. (To avoid confusion, the superscripts here marks the domain over which the model is formulated.)
Combining that with “sharpness of the phase transition”, i.e. exponential decay of correlations in the infinite-volume limit up to the threshold of symmetry breaking, we learn that for any $b>b_{c}$ the above upper bound decays exponentially, as stated in (7.3).
Thus, the kernel, which is defined in (6.11), through maximization of $\langle S^{z}_{u};S^{z}_{v}\rangle_{\sigma_{B}}^{W}$ satisfies
@@
It seems reasonable to expect that a similar area low bound on the entropy of entanglement is valid also for each of the extremal ($+$ or --) ground states throughout the phase which is characterized by symmetry breaking. However in that case while the conditional variational distance $\delta_{\boldsymbol{\sigma}_{B}}(A|C)$ may be small for configurations which are aligned with the sign of the state’s order parameter, it may not be small for some configurations $\sigma_{B}$ which are magnetised in the opposite direction. Still, with $\delta_{B}(A|C)$ being an average of the above, there is a reason to expect the criterion of Corollary 4.4 to be applicable also in that case - provided progress is made on the above point.
@@
Among the numerous other example of stoquastic Hamiltonians to which the above strategy may be of relevance is the quantum Heisenberg model [26, 50, 51, 49], its ‘flattened’ projection based interactions, studied by related methods in [42, 6, 4] (and by different methods in many other works), and also the ground states of Kitaev’s toric code [40, 49].

### v2

@@ 4.2 An improved multiscale analysis
    For ground-states of the quantum Ising model existing techniques yield $l_{0}(A)\propto\ln|\partial A|$ for general regular $A$, which results in a logarithmically corrected area law unless $d=1$.
  3. 3.
    The general results presented here have some partial overlap with theorems that were formulated for different classes of states, under different assumptions and restrictions:
    1. (a)
      Various facets of the area-law conjecture for ground states were settled for $d=1$ as well as for exactly solvable or non-interacting systems (cf. [28] for a review). In a trailblazing work, Hastings [35] proved it for ground states of one dimensional systems with gapped local Hamiltonians. The argument was streamlined in the work of Brandão and Horodecki [16], which clarified that the argument is, in essence, based on the state’s exponential clustering property.
    2. (b)
      For $d=2$, under the assumption of frustration-freeness of a gapped local Hamiltonian on $W\subset\mathbb{Z}^{2}$, Anshu, Arad and Vidick [9] devised a technique, based on a so-called detectability argument, for constructing an approximate ground state projector. Using it, Anshu, Arad and Gosset [8] established an area-law bound with a poly-logarithmic correction for the entanglement entropy of slub subsets $A$ of rectangular sets $W$.
    3. (c)
      Beyond the above two-dimensional case, only conditional results with assumptions that do not cover the full non-critical regime are available [43, 22, 15]. The area law is also known to fail for some non-regular graphs [2].
  4. 4.
    An alternative measure of entanglement is provided by the Rényi entropy $S_{\alpha}(\varrho):=(1-\alpha)^{-1}\ln\mathrm{tr}\varrho^{\alpha}$. Since it is monotone decreasing in $\alpha\in(0,\infty)$ and converges to the von Neumann entropy for $\alpha\to 1$, the above-discussed area-law bounds extend to all values $\alpha\geq 1$. In particular, that applies to integer values $\alpha\in\{2,3,\dots\}$.
  5. 5.
    For $d=1$ and nearly-critical ground states, Cardy and Calabrese [20] predicted the asymptotic value of the Rényi $2$ entanglement entropy, expressed in terms of the correlation length and the central charge of the conformal field theory describing the critical ground state.
@@ 4.3 Entropy approximations
$\frac{1}{2}\Delta_{l}$ $:=\frac{1}{2}\sum_{j=1}^{\nu^{|B_{l}|}}\left|\lambda_{j}^{\downarrow}-\mu_{j}^{\downarrow}\right|\leq\frac{1}{2}\left\|\varrho_{A}-\widehat{\varrho}_{A}\right\|_{1}\leq\sqrt{2(1-\operatorname{Re\,}\langle\psi|\psi_{(B_{l})}\rangle}$
@@
where we used [12, Lemma IV.3.2], the variational characterization of the trace norm, $\|a\|_{1}=\sup_{\|b\|=1}\left|\mathrm{tr}(ab)\right|$, as well as (3). The last line is from the fidelity bound (3.6) and the assumed decay. The Fannes-type entropy bound from Proposition A.1 thus yields
@@
In the subsequent analysis of the mutual information, we need a minor modification of Theorem 4.8 involving the approximation of $\psi$ by the vector $|\psi_{(B_{l})}\rangle$ from (2.9) with phases now optimized according to (3.4). For simplicity, we restrict attention to the case $n=2$, i.e., the phases will be adapted to the decomposition of $A$ into two separated components $A_{1},A_{2}\subset W$, cf. Figure 2. To formulate a bound, we require the following modification of Defintion 4.5.
@@
Remark:  It will be shown in Theorem 6.1 that in the above situation
@@
The proof starts by noting that for any of the three reduced states $\varrho_{D}\equiv\varrho_{D}(\psi)$ with $D\in\{A,A_{1},A_{2}\}$, the arguments in the derivation of (4.9) still yield for any $l\geq l_{0}(A)$:
@@ 5 Mutual information bounds
Let $|\psi\rangle$ be a state, which is strongly conditionally exponentially decoupling beyond $l_{0}(A)=c\ln|\partial A|$ for some $c\in(0,\infty)$, and $A=A_{1}\sqcup A_{2}$ with $A_{1},A_{2}\subset W$ non-empty squares, which are separated by distance $d(A_{1},A_{2})\geq 3l_{0}(A)$. Then there are constants $C,\kappa,\eta\in(0,\infty)$, which are independent of $A_{1},A_{2}$, such that
@@
Moreover, as is evident from the subsequent proof, Theorem 5.1 admits a generalization to the mutual information of finitely many disjoint squares $A_{1},\dots,A_{n}$.
@@
    $\left|1-\langle\psi^{(12)}_{(B)}|\psi_{(B)}\rangle\right|$ $\leq\sum_{\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B}}\Big{|}\ p(\boldsymbol{\sigma}_{B})p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B})-\prod_{j=1}^{2}p(\boldsymbol{\sigma}_{B_{j}})p(\boldsymbol{\sigma}_{A_{j}}|\boldsymbol{\sigma}_{B_{j}})\Big{|}$
@@
The trace distances occurring in the application of Proposition A.1 are estimated by fidelities using (3), which is upper bounded using (2):
$2\left(1-\operatorname{Re\,}\langle\psi^{(12)}_{(B)}|\psi_{(B)}\rangle\right)$ $\leq 2\delta\left(A_{1}\sqcup B_{1}|A_{2}\sqcup B_{2}\right)$
$\leq 2e^{-(l-l_{0}(A))/(3\xi)}=2\ |\partial A|^{c/(3\xi)}e^{-l/(3\xi)}\,.$ (5.6)
The last follows from the assumption and the fact that $d(A_{1}\sqcup B_{1},A_{2}\sqcup B_{2})\geq l/3$ by construction. Since $|\partial A|\leq 2\max\left\{|\partial A_{1}|,|\partial A_{2}|\right\}$, this completes the proof. ∎
@@ 6.1 General properties
${\delta}(A\,|\,C)=\frac{1}{2}\max_{\|F\|_{\infty}\leq 1}\sum_{\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{C}}F(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{C})\left[p(\boldsymbol{\sigma}_{A}\boldsymbol{\sigma}_{C})-p(\boldsymbol{\sigma}_{A})p(\boldsymbol{\sigma}_{C}))\right]\,,$ (6.2)
where the maximum is over $F:\Omega_{A}\times\Omega_{C}\to[-1,1]$ and attained at $F(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{C}):=\operatorname{sign}[p(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{C})-p(\boldsymbol{\sigma}_{A})p(\boldsymbol{\sigma}_{C})\big{]}$. Monotonicity in $C$ follows, since replacing $\boldsymbol{\sigma}_{C}$ by $\boldsymbol{\sigma}_{C\sqcup D}$ only enlarges the collection of functions over which the maximization is performed.
@@
${\delta}_{B}(A|\{u\})\leq\sum_{\boldsymbol{\sigma}_{B}}p(\boldsymbol{\sigma}_{B})\ \overline{{\mathrm{TV}}}_{\boldsymbol{\sigma}_{B}}(A|\{u\})\,.$ (6.6)
@@ 6.2 Relation to correlation function using the FKG property
    and $\langle\sigma_{a};\sigma_{u}\rangle_{\boldsymbol{\sigma}_{B}}:=\langle\sigma_{a}\sigma_{u}\rangle_{\boldsymbol{\sigma}_{B}}-\langle\sigma_{a}\rangle_{\boldsymbol{\sigma}_{B}}\langle\sigma_{u}\rangle_{\boldsymbol{\sigma}_{B}}$ with the subscript indicating that the expectation values are with respect to the conditional probability $p(\cdot|\boldsymbol{\sigma}_{B})$, conditioned on the value of $\boldsymbol{\sigma}_{B}$.
@@
$=\min\left\{\mathbb{P}\left(\boldsymbol{\sigma}_{A}^{(1)}\neq\boldsymbol{\sigma}_{A}^{(2)}\right)\ \big{|}\ \mbox{couplings with $\boldsymbol{\sigma}_{A}^{(1)}\rightsquigarrow\boldsymbol{\sigma}_{A}^{+},\boldsymbol{\sigma}_{A}^{(2)}\rightsquigarrow\boldsymbol{\sigma}_{A}^{-}$}\right\}\,.$
@@
$\mathbb{P}\big{(}\boldsymbol{\sigma}_{A}^{(1)}\neq\boldsymbol{\sigma}_{A}^{(2)}\big{)}$ $\leq\mathbbm{E}\left(\sum_{a\in A}(\sigma_{a}^{(1)}-\sigma_{a}^{(2)})/2\right)$
$=\frac{1}{2}\sum_{\boldsymbol{\sigma}_{A}}M_{A}(\boldsymbol{\sigma}_{A})\big{[}p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B}\sigma_{u}=+1)-p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B}\sigma_{u}=-1)\big{]}$
@@
$=\frac{1}{2}\sum_{\sigma_{u},\widehat{\sigma}_{u}}\sum_{\boldsymbol{\sigma}_{A},\widehat{\boldsymbol{\sigma}}_{A}}p(\boldsymbol{\sigma}_{A},\sigma_{u}|\boldsymbol{\sigma}_{B})\ p(\widehat{\boldsymbol{\sigma}}_{A},\widehat{\sigma}_{u}|\boldsymbol{\sigma}_{B})\left[M_{A}(\boldsymbol{\sigma}_{A})-M_{A}(\widehat{\boldsymbol{\sigma}}_{A})\right]\left[\sigma_{u}-\widehat{\sigma}_{u}\right]$
$=2\,p_{\boldsymbol{\sigma}_{B}}(+)\ p_{\boldsymbol{\sigma}_{B}}(-)\sum_{\boldsymbol{\sigma}_{A}}M_{A}(\boldsymbol{\sigma}_{A})\big{[}p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B}\sigma_{u}=+1)-p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B}\sigma_{u}=-1)\big{]}$ (6.13)
@@ 7.1 Explicit results
    FKG: Like its classical version [38, 32], the quantum Ising model’s probabilty has the FKG property, in the natural partial order of the stoquastic $S^{z}$ basis (cf. [33]). This makes Theorem 6.4 applicable to this model, with its variables $\sigma_{u}$ interpreted as $S^{z}_{u}$.
@@
    In its classical version, the inequality is a direct implication of the recently formulated Ising inequality, which was presented in [24, Eq. (1.4)]. The validity of its quantum version follows by a weak continuity argument, similar to that spelled in the discussion of the FKG inequality for QIM in [33]. An alternative direct proof, of both classical and quantum versions using the random current approach, is included in [3].
@@
The ground state of the quantum Ising model in a finite volume $W\subset\mathbb{Z}^{d}$ with free boundary conditions is a unique pure state at $b>0$ due to the Perron-Frobenius theorem and stoquastic in the $z$-diagonalising computational basis. By the extended Ding-Song-Sun inequality (7.4), for any finite volume and throughout the range of the model’s coupling constants, its ground state’s truncated correlation function, conditioned on arbitrary configuration of the buffer, satisfies:
@@
where the last upper bound holds thanks to the Griffiths inequality.
Combining that with “sharpness of the phase transition”, i.e. exponential decay of correlations in the infinite-volume limit up to the threshold of symmetry breaking, we learn that for any $b>b_{c}$ the above upper bound decays exponentially, as stated in (7.3). Thus, the kernel, which is defined in (6.11), through maximization of $\langle S^{z}_{u};S^{z}_{v}\rangle_{\sigma_{B}}^{(W)}$ satisfies
@@ 7.2 Discussion
  1. 1.
    In previous studies of the entanglement in QIM [33, 21, 34] decorrelation effects were expressed through a ratio weak-mixing estimate, in which the proximity of two probability distributions $p(\boldsymbol{\sigma}_{A})$ and $\widehat{p}(\boldsymbol{\sigma}_{A})$ is approached through an upper bound on $|p(\boldsymbol{\sigma}_{A})/\widehat{p}(\boldsymbol{\sigma}_{A})-1|$. This approach works well when $|\delta p|(\boldsymbol{\sigma}_{A}):=|p(\boldsymbol{\sigma}_{A})-\widehat{p}(\boldsymbol{\sigma}_{A})|$ is typically much smaller than $p(\boldsymbol{\sigma}_{A})$ itself. However that condition fails in dimensions $d>1$ even if $\delta p$ is of order $e^{-L(A)/\xi}$, since there $p(\boldsymbol{\sigma}_{A})$ is of the order $e^{-\alpha L(A)^{d}}$. In one dimension it applies but only where the correlation length is shorter than a finite threshold. This limitation is avoided in our estimate of the trace distance between the corresponding states through a more effective use of the FKG inequality.
  2. 2.
    It seems reasonable to expect that an entanglement entropy bound similar to what is proven here for subcritical states is valid also for each of the pure ground states of broken symmetry, away from the critical case. In that case the conditional variational distance $\delta_{\boldsymbol{\sigma}_{B}}(A|C)$ may be small for configurations of the buffer that are aligned with the state’s order parameter, but not so for configurations $\boldsymbol{\sigma}_{B}$ which are magnetised in the opposite direction. Still, with $\delta_{B}(A|C)$ being an average of the above, one may expect the criterion of Corollary 4.7 to be applicable also in that case - provided progress is made on a more effective handling of the above observations.
@@
  3. 3.
    Among the numerous other examples of stoquastic Hamiltonians to which the above strategy may be of relevance is the quantum Heisenberg model [27, 50, 51, 49], its ‘flattened’ projection based interactions, studied by related methods in [42, 6, 4] (and by different methods in many other works), and also the ground states of Kitaev’s toric code [40, 49].